Consider a local utility that has a demand curve given by p = 100,000 - 2q and has constant marginal cost of $2. The fixed cost of the firm is $100,000.

a. What is the Average total cost (ATC) of the firm?

b. Find the unregulated monopoly price and output.

c. Calculate the deadweight loss or the welfare loss.

d. Where is the efficient output?

e. Find the price that needs to be set to regulate it using the average cost pricing rule.
Answer #1

P=10oo0o-29 mc=2 fe-loooo © TE= Flare = 100000-422 ATC= 10000MG - 72 Profit is max when MR-me 100000-49 = 2 q- 999998 * 24999